OpenStudy (star):

find the particular solution of y'' + y' = x + e^x satisfying y(0)=0 and y'(0)=1

myininaya (myininaya):

\[y=Ax^2+Bx+Cxe^x+De^x\] \[y'=2Ax+B+Ce^x+Cxe^x+De^x\] \[y''=2A+Ce^x+Ce^x+Cxe^x+De^x\] so we have y''+y'= \[[xe^x(C)+e^x(2C+D)+2A]+[xe^x(C)+e^x(C+D)+x(2A)+B]\] \[=xe^x(2C)+e^{x}(3C+2D)+x(2A)+(2A+B)\] this is suppose to equal x+e^x so we have \[2C=0; 3C+2D=1; 2A=1; 2A+B=0\] \[C=0;D=\frac{1}{2}; A=\frac{1}{2}; B=-1\]

myininaya (myininaya):

\[y=\frac{1}{2}x^2-x+\frac{1}{2}e^x\]

myininaya (myininaya):

+constant

myininaya (myininaya):

you also need to find the homogeneous solution

myininaya (myininaya):

\[y''+y'=0\] solve this

myininaya (myininaya):

\[r^2+r=0 => r(r+1)=0 => r=0; r=-1=>y_h=c_1+c_2e^{-t}\]

myininaya (myininaya):

\[y=\frac{1}{2}x^2-x+\frac{1}{2}e^x+c_2e^{-t}+c\]

myininaya (myininaya):

oops t is x

myininaya (myininaya):

now you need to apply your initial conditions